OpenStudy (andijo76):

if np >5 and nq>5 estimate P (fewer than 4) with n=13 p=0.4 by using normal distribution as an approximate to the binomial distribution if np <5 or nq< then state that the normal approximation is not suitable

OpenStudy (andijo76):

q=1-p -1-0.4 -0.6 np-13*0.4=5.2 nq= 13*0.6= 7.8 np=5.2 nq=7.8 μ=np =13*0.4 =5.2 \[\sqrt{13∗0.4∗0.6}\]=1.766 x=5-0.5 =4.5 x-u/o =(5-0.5)/-6=1.766 =0.8493

OpenStudy (andijo76):

there is one more part but i cant figure it out HELP